# Mastering AI: Big Data, Deep Learning, and the Evolution of Large Language Models - Blockchain and Applications

### v2

@@ 8.3.2 Definition and Examples of Hard Forks
A well-known example of a hard fork is the creation of Bitcoin Cash (BCH) in 2017[180]. This fork was driven by disagreements within the Bitcoin community about how to scale the network, particularly regarding the block size limit. Bitcoin Cash increased the block size to allow more transactions per block, whereas Bitcoin (BTC) retained its smaller block size. The two networks became entirely separate, with their own coins and ecosystems.
Another example of a hard fork is the Ethereum DAO fork in 2016[184]. After a vulnerability in the DAO (Decentralized Autonomous Organization) contract was exploited, a large portion of Ethereum’s funds was stolen. The community decided to implement a hard fork to reverse the effects of the hack, creating two separate blockchains: Ethereum (ETH)[185] and Ethereum Classic (ETC)[186].
@@ 8.3.3 SegWit Upgrade and Segregated Witness
Segregated Witness, commonly known as SegWit[183], was a major Bitcoin upgrade implemented through a soft fork in 2017. It aimed to solve a critical problem known as transaction malleability, which made it possible for attackers to change small parts of a transaction’s identifier without altering its content. This issue made it difficult for users to track and verify transactions accurately.
@@ 8.4.1 The Birth of Bitcoin Cash and Its Causes
Bitcoin Cash (BCH) was created in August 2017[180] as a result of a hard fork from Bitcoin. The primary reason for the fork was a long-standing debate within the Bitcoin community over how to best scale the network to accommodate increasing transaction demand. Bitcoin’s block size, capped at 1MB, limited the number of transactions that could be processed in each block, leading to higher fees and slower transaction times during periods of high demand.
One side of the debate argued that increasing the block size was the simplest and most effective solution to improve Bitcoin’s scalability[173]. However, another faction, including Bitcoin Core developers, opposed increasing the block size, citing concerns about centralization (larger blocks require more storage and processing power) and the potential risks of altering Bitcoin’s fundamental design[187].
@@ 8.4.2 Technical Changes in Bitcoin Cash: 8MB Block Size
The primary technical change introduced by Bitcoin Cash was the increase of the block size limit from 1MB to 8MB[188]. By allowing blocks to be up to 8 times larger, Bitcoin Cash aimed to process more transactions in each block, thereby reducing transaction confirmation times and lowering transaction fees during periods of high network activity.
@@ 8.5.1 Layer 2 Solutions
Layer 2 solutions are protocols built on top of the base blockchain (Layer 1) to improve scalability and reduce fees[189]. Instead of changing the core blockchain protocol, these solutions create an additional layer where transactions can occur off-chain, reducing the load on the base layer.
@@ 9.1.1 Public Key Cryptography and Digital Signature Assurance
Bitcoin relies heavily on public key cryptography to secure its transactions and ensure the integrity and ownership of the funds[190]. Each Bitcoin user has a pair of cryptographic keys: a public key and a private key. The public key can be shared with anyone, and it serves as an address to receive Bitcoin. However, the private key must remain confidential, as it is used to sign transactions and authorize the transfer of Bitcoin.
When a user wants to send Bitcoin, they create a transaction message that includes details such as the amount of Bitcoin and the recipient’s public address[191]. This message is then signed using the sender’s private key. This digital signature is crucial because it proves that the transaction was indeed initiated by the rightful owner of the Bitcoin without revealing the private key itself[192]. The signature is verified using the corresponding public key, ensuring that the transaction is authentic and not tampered with.
@@
To illustrate this process in Python, we can simulate public key cryptography using the ‘ecdsa‘ (Elliptic Curve Digital Signature Algorithm) library[193]. This simple example demonstrates generating a private-public key pair, signing a message, and verifying the signature.
@@ 9.1.2 Hash Locking and Time Locking
Hash locking ensures that certain conditions are met before funds can be transferred[194]. It works by locking the Bitcoin transaction with a specific cryptographic condition, typically a hash pre-image. A hash lock requires that the correct input (pre-image) be provided to unlock the funds. This mechanism is often used in ”Atomic Swaps”[195] or ”Hashed Time-Locked Contracts (HTLCs)”[196] in the context of the Lightning Network, where two parties exchange Bitcoin without the need for a third-party intermediary.
@@
Time locking is another mechanism that delays the spending of a transaction until a specified time or block height[197]. Bitcoin has two types of time locks:
  • CheckLockTimeVerify (CLTV): Prevents a transaction from being spent until a specified block height or time[198].
  • CheckSequenceVerify (CSV): Introduces relative time locking, where a transaction can only be spent after a certain number of blocks have passed[198].
@@ 9.2.1 Double-Spending Attack
A double-spending attack occurs when a malicious actor attempts to spend the same Bitcoin more than once[199]. Bitcoin’s decentralized nature and the delay between transaction broadcasts create a potential window for such an attack. If an attacker can convince one party to accept a transaction and then spends the same Bitcoin elsewhere before the network has confirmed the original transaction, they have effectively double-spent.
Bitcoin mitigates this risk using its Proof of Work (PoW) consensus mechanism[149], where miners expend computational power to validate transactions and include them in blocks. Once a block is added to the blockchain, it becomes increasingly difficult to alter or reverse the transactions within that block. As more blocks are added on top, the security of those transactions increases.
@@
The probability of a successful double-spend decreases significantly after multiple confirmations. Generally, six confirmations are considered safe, meaning the transaction is buried under six additional blocks[160].
@@ 9.2.2 Principle and Impact of a 51% Attack
A 51% attack refers to a situation where a single entity or group of miners gains control of more than 50% of the network’s total mining hash rate[200]. In such a scenario, the attacker could potentially: - Reverse recent transactions, allowing them to double-spend Bitcoin. - Prevent new transactions from gaining confirmations. - Halt the network temporarily by preventing new blocks from being mined.
@@ 9.3.1 Pseudonymity and User Privacy
Bitcoin operates on a pseudonymous system, where users are not required to reveal their real identities. Instead, they are identified by unique addresses, which are derived from cryptographic keys[201]. Each Bitcoin address is a long string of alphanumeric characters, and it serves as a pseudonym for the user.
@@ 9.3.2 Transaction Traceability and Coin Mixing Services
To address these concerns, several privacy-enhancing tools have emerged. One of the most popular methods is the use of coin mixing services, such as CoinJoin[202]. These services help obfuscate the trail of Bitcoin transactions by combining the transactions of multiple users into a single large transaction. This makes it difficult to trace the flow of funds from one user to another.
@@ 9.4.1 Bitcoin Improvement Proposal (BIP) Process
Bitcoin’s development and improvement are guided by a structured process known as the Bitcoin Improvement Proposal (BIP)[203]. The BIP process allows developers to propose new features, enhancements, or changes to the Bitcoin protocol in an organized and transparent manner. This system ensures that changes to Bitcoin are thoroughly vetted and reviewed by the community before they are implemented.
@@
Example: One of the most famous BIPs is BIP 141[183], which proposed the implementation of Segregated Witness (SegWit). SegWit introduced changes to the way transactions were structured, increasing the capacity of Bitcoin’s blocks and solving some issues related to transaction malleability. The proposal underwent extensive community discussion and testing before being adopted by the network.
@@ 9.4.2 Decentralized Governance of the Bitcoin Community
Example: The 2017 block size debate highlighted how Bitcoin’s decentralized governance works[204]. The community was divided on whether to increase the block size to allow for more transactions per block. Some developers and miners supported the change, while others opposed it, leading to a hard fork that resulted in the creation of Bitcoin Cash. This event showcased how the decentralized governance model allows for differences of opinion, with each faction free to pursue its vision of Bitcoin’s future.
@@ 10.1.1 Vitalik Buterin’s Vision
Vitalik Buterin, a programmer and visionary in the blockchain space, created Ethereum in 2015[205]. His journey towards developing Ethereum began in 2013 when he recognized the limitations of Bitcoin, the first and most popular cryptocurrency at the time. Bitcoin was primarily designed as a peer-to-peer electronic cash system, allowing for secure and decentralized money transfers. However, it was limited in its functionality, as its blockchain could only process simple transactions and had no ability to support complex applications.
Buterin saw an opportunity to expand on Bitcoin’s concept by creating a new blockchain that could serve as a platform for not just cryptocurrency transactions, but also decentralized applications (dApps)[205]. His vision was to build a system where developers could create applications that operate autonomously on a blockchain, with no reliance on intermediaries. These applications would use ”smart contracts,” which are self-executing contracts with the terms of the agreement directly written into code.
@@ 10.1.2 Turing Completeness and Ethereum’s Design Goals
A key element of Ethereum’s innovation is its Turing-complete programming environment[206]. Turing completeness refers to a system’s ability to perform any calculation or solve any problem that can be computed, provided enough time and resources are available. In simpler terms, a Turing-complete system can execute a full range of computational instructions[206].
@@ 10.2.1 Smart Contracts and Decentralized Applications (dApps)
Smart Contracts: A smart contract is a self-executing piece of code with the terms of the agreement written directly into the code[207]. When certain predefined conditions are met, the contract automatically executes without the need for intermediaries. For example, a smart contract could be written to automatically transfer ownership of a digital asset once payment is received. This automation eliminates the need for third parties like banks or escrow services.
@@
Decentralized Applications (dApps): Developers can build decentralized applications (dApps) on Ethereum[208], which operate without centralized servers. These dApps are powered by smart contracts, enabling them to function automatically based on the code written into the blockchain. This allows for a wide range of applications, from decentralized finance (DeFi) platforms to games and social networks, all running without a central authority.
@@ 10.2.2 Role of the Ethereum Virtual Machine (EVM)
At the heart of Ethereum’s functionality is the Ethereum Virtual Machine (EVM)[209]. The EVM is a global, decentralized computer that enables developers to execute smart contracts on the Ethereum blockchain. It provides a runtime environment for smart contracts, ensuring that they are executed consistently across all nodes in the network.
@@ 10.2.3 Ethereum’s Gas Mechanism
One of Ethereum’s critical innovations is the concept of ”Gas.”[210] Gas is a unit that measures the computational work required to execute transactions and smart contracts on the Ethereum network.
@@
[…]
@@
The Gas mechanism is essential for maintaining the stability, security, and efficiency of the Ethereum network. By attaching a cost to computational work, Ethereum ensures that resources are used judiciously and that the network remains operational even as it scales[211].
@@ 11.1 Definition of Smart Contracts
Smart contracts are digital agreements that are automatically enforced by code. Unlike traditional contracts, which depend on third parties (such as lawyers or courts) for enforcement, smart contracts are self-executing. This concept was first introduced by Nick Szabo in the 1990s[207].
@@ 11.1.1 Nick Szabo’s Concept of Smart Contracts
Nick Szabo is a computer scientist, legal scholar, and cryptographer. In the 1990s, Szabo proposed the idea of smart contracts to improve the way agreements are made and enforced[212]. He described smart contracts as a set of promises, specified in digital form, including protocols within which the parties perform on these promises.
The core idea behind Szabo’s concept was that contractual clauses can be embedded in hardware and software systems, making them self-enforcing[213]. For example, in a vending machine, a customer inserts coins and selects a product, and the machine automatically delivers the product when the correct amount is inserted. In this analogy, the machine acts as a smart contract, enforcing the terms (payment and delivery) without the need for a human intermediary. This concept has been greatly expanded with the advent of blockchain technology.
@@ 11.2.1 Writing and Deploying Smart Contracts
On Ethereum, smart contracts are written using a programming language called Solidity[214]. Developers create smart contracts by writing code that specifies the contract’s terms and conditions. Once the code is written, it is compiled and deployed to the Ethereum blockchain.
@@ 11.2.2 Solidity Programming Language
Solidity is the most widely-used language for writing smart contracts on Ethereum[185]. It is a statically-typed language that is designed to work seamlessly with the Ethereum Virtual Machine (EVM). Solidity features a syntax similar to JavaScript, which makes it approachable for developers familiar with web development.

### v3

@@ 8.3.2 Definition and Examples of Hard Forks
A well-known example of a hard fork is the creation of Bitcoin Cash (BCH) in 2017[Aggarwal2019ASA]. This fork was driven by disagreements within the Bitcoin community about how to scale the network, particularly regarding the block size limit. Bitcoin Cash increased the block size to allow more transactions per block, whereas Bitcoin (BTC) retained its smaller block size. The two networks became entirely separate, with their own coins and ecosystems.
Another example of a hard fork is the Ethereum DAO fork in 2016[Mehar2021UnderstandingAR]. After a vulnerability in the DAO (Decentralized Autonomous Organization) contract was exploited, a large portion of Ethereum’s funds was stolen. The community decided to implement a hard fork to reverse the effects of the hack, creating two separate blockchains: Ethereum (ETH)[ethereumCompleteGuide] and Ethereum Classic (ETC)[ethereumclassicEthereumClassic].
@@ 8.3.3 SegWit Upgrade and Segregated Witness
Segregated Witness, commonly known as SegWit[bitcoin0141Bitcoin], was a major Bitcoin upgrade implemented through a soft fork in 2017. It aimed to solve a critical problem known as transaction malleability, which made it possible for attackers to change small parts of a transaction’s identifier without altering its content. This issue made it difficult for users to track and verify transactions accurately.
@@ 8.4.1 The Birth of Bitcoin Cash and Its Causes
Bitcoin Cash (BCH) was created in August 2017[Aggarwal2019ASA] as a result of a hard fork from Bitcoin. The primary reason for the fork was a long-standing debate within the Bitcoin community over how to best scale the network to accommodate increasing transaction demand. Bitcoin’s block size, capped at 1MB, limited the number of transactions that could be processed in each block, leading to higher fees and slower transaction times during periods of high demand.
One side of the debate argued that increasing the block size was the simplest and most effective solution to improve Bitcoin’s scalability[GroupBlockSI]. However, another faction, including Bitcoin Core developers, opposed increasing the block size, citing concerns about centralization (larger blocks require more storage and processing power) and the potential risks of altering Bitcoin’s fundamental design[deFilippi2016TheIP].
@@ 8.4.2 Technical Changes in Bitcoin Cash: 8MB Block Size
The primary technical change introduced by Bitcoin Cash was the increase of the block size limit from 1MB to 8MB[Johnson2021DoesBC]. By allowing blocks to be up to 8 times larger, Bitcoin Cash aimed to process more transactions in each block, thereby reducing transaction confirmation times and lowering transaction fees during periods of high network activity.
@@ 8.5.1 Layer 2 Solutions
Layer 2 solutions are protocols built on top of the base blockchain (Layer 1) to improve scalability and reduce fees[Mandal2023InvestigatingLS]. Instead of changing the core blockchain protocol, these solutions create an additional layer where transactions can occur off-chain, reducing the load on the base layer.
@@ 9.1.1 Public Key Cryptography and Digital Signature Assurance
Bitcoin relies heavily on public key cryptography to secure its transactions and ensure the integrity and ownership of the funds[Conti2017ASO]. Each Bitcoin user has a pair of cryptographic keys: a public key and a private key. The public key can be shared with anyone, and it serves as an address to receive Bitcoin. However, the private key must remain confidential, as it is used to sign transactions and authorize the transfer of Bitcoin.
When a user wants to send Bitcoin, they create a transaction message that includes details such as the amount of Bitcoin and the recipient’s public address[Puranam2019AnatomyAL]. This message is then signed using the sender’s private key. This digital signature is crucial because it proves that the transaction was indeed initiated by the rightful owner of the Bitcoin without revealing the private key itself[Vesel2021BitcoinAQ]. The signature is verified using the corresponding public key, ensuring that the transaction is authentic and not tampered with.
@@
To illustrate this process in Python, we can simulate public key cryptography using the ‘ecdsa‘ (Elliptic Curve Digital Signature Algorithm) library[Johnson1998EllipticCD]. This simple example demonstrates generating a private-public key pair, signing a message, and verifying the signature.
@@ 9.1.2 Hash Locking and Time Locking
Hash locking ensures that certain conditions are met before funds can be transferred[Churchill2015WhySW]. It works by locking the Bitcoin transaction with a specific cryptographic condition, typically a hash pre-image. A hash lock requires that the correct input (pre-image) be provided to unlock the funds. This mechanism is often used in ”Atomic Swaps”[Miraz2019AtomicCS] or ”Hashed Time-Locked Contracts (HTLCs)”[Bribing2021TimelockedB] in the context of the Lightning Network, where two parties exchange Bitcoin without the need for a third-party intermediary.
@@
Time locking is another mechanism that delays the spending of a transaction until a specified time or block height[PrasannaSrinivasan2021NetworkVI]. Bitcoin has two types of time locks:
  • CheckLockTimeVerify (CLTV): Prevents a transaction from being spent until a specified block height or time[lightningTimelocksBuilders].
  • CheckSequenceVerify (CSV): Introduces relative time locking, where a transaction can only be spent after a certain number of blocks have passed[lightningTimelocksBuilders].
@@ 9.2.1 Double-Spending Attack
A double-spending attack occurs when a malicious actor attempts to spend the same Bitcoin more than once[Kumar2023ARO]. Bitcoin’s decentralized nature and the delay between transaction broadcasts create a potential window for such an attack. If an attacker can convince one party to accept a transaction and then spends the same Bitcoin elsewhere before the network has confirmed the original transaction, they have effectively double-spent.
Bitcoin mitigates this risk using its Proof of Work (PoW) consensus mechanism[ferdous2020blockchainconsensusalgorithmssurvey], where miners expend computational power to validate transactions and include them in blocks. Once a block is added to the blockchain, it becomes increasingly difficult to alter or reverse the transactions within that block. As more blocks are added on top, the security of those transactions increases.
@@
The probability of a successful double-spend decreases significantly after multiple confirmations. Generally, six confirmations are considered safe, meaning the transaction is buried under six additional blocks[Wan2019EvaluatingTI].
@@ 9.2.2 Principle and Impact of a 51% Attack
A 51% attack refers to a situation where a single entity or group of miners gains control of more than 50% of the network’s total mining hash rate[AponteNovoa2021The5A]. In such a scenario, the attacker could potentially: - Reverse recent transactions, allowing them to double-spend Bitcoin. - Prevent new transactions from gaining confirmations. - Halt the network temporarily by preventing new blocks from being mined.
@@ 9.3.1 Pseudonymity and User Privacy
Bitcoin operates on a pseudonymous system, where users are not required to reveal their real identities. Instead, they are identified by unique addresses, which are derived from cryptographic keys[Chang2020ImprovingBO]. Each Bitcoin address is a long string of alphanumeric characters, and it serves as a pseudonym for the user.
@@ 9.3.2 Transaction Traceability and Coin Mixing Services
To address these concerns, several privacy-enhancing tools have emerged. One of the most popular methods is the use of coin mixing services, such as CoinJoin[coinjoinCoinJoinBitcoin]. These services help obfuscate the trail of Bitcoin transactions by combining the transactions of multiple users into a single large transaction. This makes it difficult to trace the flow of funds from one user to another.
@@ 9.4.1 Bitcoin Improvement Proposal (BIP) Process
Bitcoin’s development and improvement are guided by a structured process known as the Bitcoin Improvement Proposal (BIP)[githubGitHubBitcoinbips]. The BIP process allows developers to propose new features, enhancements, or changes to the Bitcoin protocol in an organized and transparent manner. This system ensures that changes to Bitcoin are thoroughly vetted and reviewed by the community before they are implemented.
@@
Example: One of the most famous BIPs is BIP 141[bitcoin0141Bitcoin], which proposed the implementation of Segregated Witness (SegWit). SegWit introduced changes to the way transactions were structured, increasing the capacity of Bitcoin’s blocks and solving some issues related to transaction malleability. The proposal underwent extensive community discussion and testing before being adopted by the network.
@@ 9.4.2 Decentralized Governance of the Bitcoin Community
Example: The 2017 block size debate highlighted how Bitcoin’s decentralized governance works[Atzori2017BlockchainTA]. The community was divided on whether to increase the block size to allow for more transactions per block. Some developers and miners supported the change, while others opposed it, leading to a hard fork that resulted in the creation of Bitcoin Cash. This event showcased how the decentralized governance model allows for differences of opinion, with each faction free to pursue its vision of Bitcoin’s future.
@@ 10.1.1 Vitalik Buterin’s Vision
Vitalik Buterin, a programmer and visionary in the blockchain space, created Ethereum in 2015[Buterin2015ANG]. His journey towards developing Ethereum began in 2013 when he recognized the limitations of Bitcoin, the first and most popular cryptocurrency at the time. Bitcoin was primarily designed as a peer-to-peer electronic cash system, allowing for secure and decentralized money transfers. However, it was limited in its functionality, as its blockchain could only process simple transactions and had no ability to support complex applications.
Buterin saw an opportunity to expand on Bitcoin’s concept by creating a new blockchain that could serve as a platform for not just cryptocurrency transactions, but also decentralized applications (dApps)[Buterin2015ANG]. His vision was to build a system where developers could create applications that operate autonomously on a blockchain, with no reliance on intermediaries. These applications would use ”smart contracts,” which are self-executing contracts with the terms of the agreement directly written into code.
@@ 10.1.2 Turing Completeness and Ethereum’s Design Goals
A key element of Ethereum’s innovation is its Turing-complete programming environment[Tikhomirov2017EthereumSO]. Turing completeness refers to a system’s ability to perform any calculation or solve any problem that can be computed, provided enough time and resources are available. In simpler terms, a Turing-complete system can execute a full range of computational instructions[Tikhomirov2017EthereumSO].
@@ 10.2.1 Smart Contracts and Decentralized Applications (dApps)
Smart Contracts: A smart contract is a self-executing piece of code with the terms of the agreement written directly into the code[Ene2020SmartC]. When certain predefined conditions are met, the contract automatically executes without the need for intermediaries. For example, a smart contract could be written to automatically transfer ownership of a digital asset once payment is received. This automation eliminates the need for third parties like banks or escrow services.
@@
Decentralized Applications (dApps): Developers can build decentralized applications (dApps) on Ethereum[Ta2019BuildingAD], which operate without centralized servers. These dApps are powered by smart contracts, enabling them to function automatically based on the code written into the blockchain. This allows for a wide range of applications, from decentralized finance (DeFi) platforms to games and social networks, all running without a central authority.
@@ 10.2.2 Role of the Ethereum Virtual Machine (EVM)
At the heart of Ethereum’s functionality is the Ethereum Virtual Machine (EVM)[Xu2018BuildingAE]. The EVM is a global, decentralized computer that enables developers to execute smart contracts on the Ethereum blockchain. It provides a runtime environment for smart contracts, ensuring that they are executed consistently across all nodes in the network.
@@ 10.2.3 Ethereum’s Gas Mechanism
One of Ethereum’s critical innovations is the concept of ”Gas.”[Albert2018RunningOF] Gas is a unit that measures the computational work required to execute transactions and smart contracts on the Ethereum network.
@@
$\text{Total Gas Cost}=\text{Gas Limit}\times\text{Gas Price}$ (10.1)
@@
The Gas mechanism is essential for maintaining the stability, security, and efficiency of the Ethereum network. By attaching a cost to computational work, Ethereum ensures that resources are used judiciously and that the network remains operational even as it scales[Yang2019EmpiricallyAE].
@@ 11.1 Definition of Smart Contracts
Smart contracts are digital agreements that are automatically enforced by code. Unlike traditional contracts, which depend on third parties (such as lawyers or courts) for enforcement, smart contracts are self-executing. This concept was first introduced by Nick Szabo in the 1990s[Ene2020SmartC].
@@ 11.1.1 Nick Szabo’s Concept of Smart Contracts
Nick Szabo is a computer scientist, legal scholar, and cryptographer. In the 1990s, Szabo proposed the idea of smart contracts to improve the way agreements are made and enforced[Laubscher2020SmartC]. He described smart contracts as a set of promises, specified in digital form, including protocols within which the parties perform on these promises.
The core idea behind Szabo’s concept was that contractual clauses can be embedded in hardware and software systems, making them self-enforcing[Szabo1997FormalizingAS]. For example, in a vending machine, a customer inserts coins and selects a product, and the machine automatically delivers the product when the correct amount is inserted. In this analogy, the machine acts as a smart contract, enforcing the terms (payment and delivery) without the need for a human intermediary. This concept has been greatly expanded with the advent of blockchain technology.
@@ 11.2.1 Writing and Deploying Smart Contracts
On Ethereum, smart contracts are written using a programming language called Solidity[soliditylangHomeSolidity]. Developers create smart contracts by writing code that specifies the contract’s terms and conditions. Once the code is written, it is compiled and deployed to the Ethereum blockchain.
@@ 11.2.2 Solidity Programming Language
Solidity is the most widely-used language for writing smart contracts on Ethereum[ethereumCompleteGuide]. It is a statically-typed language that is designed to work seamlessly with the Ethereum Virtual Machine (EVM). Solidity features a syntax similar to JavaScript, which makes it approachable for developers familiar with web development.